# Analytic Langlands correspondence from SoV

### v1

@@ 6.2 Proof of Theorems 3 and 4
(Theorem 3) We consider the integrals $\Theta^{(n+1,1)}_{\chi}$ defined in (LABEL:Fourier(n)) in the special cases where $j_{0}=-1$ and $E_{0}=0$. With the help of the system of differential equations stated in Proposition 4 we may show that there exists a function $\Psi_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] x_{1}&\dots&x_{n}\end{smallmatrix}\big)$ satisfying
$\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{ax_{1}-b}{cx_{1}-d}&\dots&\frac{ax_{n}-b}{cx_{n}-d}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{1}{|cx_{r}+d|^{4j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big),$
which allows us to represent $\Theta^{(n+1,1)}_{\chi}$ in the form
$\Theta^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}j_{0}&\dots&j_{n}\\[3.22916pt] x_{0}&\dots&x_{n}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big).$ (6.62)
We may next note that
$\lim_{z_{0}\to\infty}\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big)=\lim_{z_{0}\to\infty}\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{z_{0}}{x_{1}-x_{0}}&\dots&\frac{z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big)$
$\hskip 156.49014pt=\prod_{r=1}^{n}{|x_{r}-x_{0}|^{4j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{1}{x_{1}-x_{0}}&\dots&\frac{1}{x_{n}-x_{0}}\end{smallmatrix}\Big)=\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big).$
The same limit of the left side of (6.62) can be computed with the help of equation (5.56). We conclude that
$\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\Theta^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big),$ (6.63)
and therefore
$\Theta^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Theta^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big).$ (6.64)
Theorem 3 now follows from (6.64) by taking the limit $z_{n}\to\infty$, $x_{n}\to\infty$ using equations (5.52) and (5.57), leading to
$\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\;\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n-1}-z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big),$ (6.65)
which is equivalent to Theorem 3. ∎
Proof.
@@
(Theorem 4) In order to prove Theorem 4 using Theorem 3 it will help to note that the Hecke operator ${\mathsf{H}}_{z_{0}}$ can be represented in terms of the limit $j_{0}\to-1$, $x_{0}\to 0$ of the operator
@@
$\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}=\lim_{j_{0}\to-1}{\mathsf{R}}_{j_{0}}\Sigma^{(n+1,1)}_{\chi}$. (6.66)
@@
The inverse $\mathsf{SoV}$-transformation $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}$ satisfies the reflection property
@@
where $\bm{j}^{\prime\prime}=(-j_{0}-1,j_{1},\dots)$ if $\bm{j}^{\prime}=(j_{0},j_{1},\dots)$.
@@
$\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\big(\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}\big)\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=$ (6.68)
$\qquad\qquad\qquad=\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
@@
$\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\pi\,\chi(z_{0},\bar{z}_{0})\,\Sigma^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
@@
$\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}=\frac{\Gamma(2j_{n}+1)}{\Gamma(-2j_{n})}\,\frac{\chi(z_{0},\bar{z}_{0})}{\chi_{n}}\,\Sigma^{(n,1)}_{\chi}$. (6.69)
Note, on the other hand, that
$\mathsf{H}_{n}\Psi_{1}:=\lim_{z_{0}\to z_{n}}|z_{r}-z_{0}|^{2j_{r}}\mathsf{H}_{z_{0}}\Psi_{1}=\frac{\Gamma(2j_{n}+1)}{\Gamma(-2j_{n})}\mathsf{H}_{\infty}\Psi_{1},$
for all $\Psi_{1}\in{\mathcal{H}}_{1}$. Taken together, we therefore find that
$\mathsf{Q}_{z_{0}}\,\Sigma^{(n,1)}_{{\chi}}=\mathsf{H}_{z_{0}}(\mathsf{H}_{n})^{-1}\,\Sigma^{(n,1)}_{{\chi}}=\frac{1}{\chi_{n}}\chi(z_{0},\bar{z}_{0})\;\Sigma^{(n,1)}_{\chi}$.
which is the Hecke eigenvalue property satisfied by $\Sigma^{(n,1)}_{\chi}$. ∎
Acknowledgements. The question about the unitarity of the $\mathsf{SoV}$-transformation was raised by D. Kazhdan in 2023 in a discussion with P. Etingof, B. Feigin, E. Frenkel, A. Goncharov, and one of the authors (J.T.). J.T. would like to thank all participants for these discussions. We’d in particular like to thank P. Etingof for pointing out reference [GrS] to us.

### v2

@@ 5.2 Comparison of limits
Concerning the limit $z_{0}\to\infty$ of the measure of integration in (5.47) note that
$\lim_{z_{0}\to\infty}z_{0}^{-1}\hat{\rho}(\bm{k})=k_{0}=-\theta(\bm{k}),$ (5.56)
on the support of $\delta^{(2)}(k_{0}+\dots+k_{n-1})$. It follows that the integration measure simplifies as
$\lim_{z_{0}\to\infty}$ $|z_{0}|^{2j_{n}}\int\prod_{r=0}^{n-1}\frac{d^{2}k_{r}}{\pi|k_{r}|^{2j_{r}+2}}\;e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}\pi\delta^{2}(k_{0}+\dots+k_{n-1})|\hat{\rho}(\bm{k})|^{-2j_{n}}$
$=\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}}{\pi|k_{r}|^{2j_{r}+2}}\;e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}|\theta(\bm{k})|^{-2j_{n}}e^{-2\mathrm{i}\,\mathrm{Im}(\theta(\bm{k})x_{0})}$
$=\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}}{\pi|k_{r}|^{2j_{r}+2}}\;e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}\int\frac{d^{2}k_{n}}{\pi|k_{n}|^{2j_{n}+2}}\;e^{2\mathrm{i}\,\mathrm{Im}(k_{n}x_{0})}\;\pi\delta^{2}(k_{n}+\theta(\bm{k}))\,|\theta(\bm{k})|^{2}.$
It remains to note that the equations defining $y_{k}(\bm{k})$ simplify for $z_{0}\to\infty$ as
$k_{r}=\frac{\hat{\rho}(\bm{k})}{z_{r}-z_{0}}\frac{\prod_{k=0}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r,0}(z_{r}-z_{s})}=\theta(\bm{k})\frac{\prod_{k=0}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r,0}(z_{r}-z_{s})}+{\mathcal{O}}(z_{0}^{-1}),$ (5.57)
in order to identify the integral defining $\Theta^{(n,0)}_{\chi}$ in (5.54), proving (5.52).
In order to verify (5.53) let us note that the leading contributions to the limit $x_{n}\to\infty$ of the left side are coming from the region where $k_{n}\to 0$. It follows from (5.55) that there must exist an index $k\in\{0,\dots,n-3\}$ such that $y_{k}\to\infty$. The permutation symmetry of the integrand allows us to assume without loss of generality that $y_{0}\to\infty$. We may next note that
$k_{r}=(-k_{n}){(z_{r}-y_{0})}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{j})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}=y_{0}k_{n}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}+\mathcal{O}(y^{-1}_{0}).$ (5.58)
However, we should recall that the equations relating $\bm{k}$ and $\bm{y}$ follow for $k_{n}\neq 0$ from
$\sum_{r=1}^{n-1}\frac{k_{n}}{y-z_{r}}=\theta(\bm{k})(y-y_{0})\frac{\prod_{k=1}^{n-3}(y-y_{k})}{\prod_{r=1}^{n-1}(y-z_{r})},$
implying, in particular, $\theta(\bm{k})=\sum_{r=1}^{n-1}k_{r}=-k_{n}$ when $k_{n}\neq 0$, and $k_{n}y_{0}=\sum_{r=1}^{n-1}z_{r}k_{r}=\rho$ in the limit $k_{n}\to 0$, $y_{0}\to\infty$.
The integrand of the integration over $k_{n}$ simplifies for $k_{n}\to 0$, $y_{0}\to\infty$ due to the asymptotic behaviour (3.34),
$\chi(y_{0},\bar{y}_{0})=|y_{0}|^{-2j_{n}}C_{+}^{(n)}\big((1+{\mathcal{O}}(|y_{0}|^{-1})\big)+|y_{0}|^{2j_{n}+2}C_{-}^{(n)}\big(1+{\mathcal{O}}(|y_{0}|^{-1})\big).$ (5.59)
By keeping in mind that $k_{n}y_{0}=\rho$ one can carry out the integral over $k_{n}$ using (5.59). One of the resulting terms is proportional to a delta-distribution in $x_{n}$ which does not contribute to the limit $x_{n}\to\infty$. The remaining term is proportional to
$\int\frac{d^{2}k_{n}}{\pi}\;\frac{e^{2\mathrm{i}\,\mathrm{Im}(k_{n}x_{n})}}{|k_{n}|^{2j_{n}}}\bigg|\frac{\rho}{k_{n}}\bigg|^{2j_{n}+2},$
which can be computed with the help of the identity ([GeS], Appendix B1.7, equation (7))
$\frac{1}{\pi}\int\frac{d^{2}k}{|k|^{4j+2}}\;e^{2\mathrm{i}\,\mathrm{Im}(kx)}=\frac{\Gamma(-2j)}{\Gamma(2j+1)}|x|^{4j}.$ (5.60)
We thereby find that the the left side of (5.53) is equal to
$C_{-}^{(n)}\frac{\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}}{\pi}\,\frac{e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\,\pi\,\delta^{(2)}\big({\textstyle\sum_{r=1}^{n-1}}k_{r}\big)|\rho(\bm{k})|^{2(j_{n}+1)}\prod_{k=1}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
which coincides with the right side of (5.53). ∎
6 Hecke eigenvalue property
@@
We shall now prove Theorem 4 using Theorem 3. To this aim let us note that the Hecke operator ${\mathsf{H}}_{z_{0}}$ can be represented in terms of the limit $j_{0}\to-1$, $x_{0}\to 0$ of the operator
@@
$\chi_{n}\,\mathbf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}=\lim_{j_{0}\to-1}{\mathsf{R}}_{j_{0}}\Sigma^{(n+1,1)}_{\chi}$. (6.61)
@@
The inverse SoV-transformation $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}$ satisfies the reflection property
@@
where $\bm{j}^{\prime\prime}=(-j_{0}-1,j_{1},\dots,j_{n})$ if $\bm{j}^{\prime}=(j_{0},j_{1},\dots,j_{n})$.
@@
$\chi_{n}\big(\mathbf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}\big)\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] 0&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$ (6.63)
@@
$\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] 0&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\chi(z_{0},\bar{z}_{0})\,\Sigma^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
@@
$\chi_{n}\mathbf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}={\chi(z_{0},\bar{z}_{0})}\,\Sigma^{(n,1)}_{\chi}$. (6.64)
Taken together, we therefore find that
$\mathbf{Q}_{z_{0}}\,\Sigma^{(n,1)}_{{\chi}}=\mathbf{H}_{z_{0}}\mathbf{K}\,\Sigma^{(n,1)}_{{\chi}}=\frac{1}{\chi_{n}}\chi(z_{0},\bar{z}_{0})\;\Sigma^{(n,1)}_{\chi}$. (6.65)
This concludes the proof of Theorem 4. $\square$.
Remark 4.
The form of the eigenvalue in (6.65) suggests that it is natural to consider the operator ${\mathbf{H}}_{\infty}:{\mathcal{H}}_{1}\to{\mathcal{H}}_{0}$ satisfying ${\mathbf{H}}_{\infty}\Sigma^{(n,1)}_{{\chi}}=\chi_{n}\Sigma^{(n,0)}_{{\chi}}$, together with $\mathbf{R}_{z_{0}}=\mathbf{H}_{z_{0}}{\mathbf{H}}_{\infty}$. The operator $\mathbf{R}_{z_{0}}$ satisfies $\mathbf{R}_{z_{0}}\Sigma^{(n,1)}_{{\chi}}=\chi(z_{0},\bar{z}_{0})\;\Sigma^{(n,1)}_{\chi}$. One may identify $\mathbf{H}$$_{\infty}$ with one of the two operators appearing in the asymptotics of ${\mathbf{H}}_{z_{0}}$ for $z_{0}\to\infty$, see [EFK4, Section 2.16.6].
Acknowledgements. The question about the unitarity of the SOV-transformation was raised by D. Kazhdan in 2023 in a discussion with P. Etingof, B. Feigin, E. Frenkel, A. Goncharov, and one of the authors (J.T.). J.T. would like to thank all participants for these discussions. We’d in particular like to thank P. Etingof for pointing out reference [GrS] to us. We’d furthermore like to thank P. Etingof and G. Felder for very useful comments on a previous version of the draft.